1) Calculate [H3O+] in the following aqueous solution at 25 ∘C: [OH−]= 1.0×10−9 M .
2) Calculate [H3O+] in the following aqueous solution at 25 ∘C: [OH−]= 2.6×10−2 M .
3) Calculate [H3O+] in the following aqueous solution at 25 ∘C: [OH−]= 7.0×10−12 M .
1
Expert's answer
2020-04-02T08:58:09-0400

The H3O+ ion is considered to be the same as the H+ ion as it is the H+ ion joined to a water molecule. The proton cannot exist in aqueous solution, due to its positive charge it is attracted to the electrons on water molecules and the symbol H3O+ is used to represent this transfer

"[H^+][OH^-]=10^{-14}\\"


1) "[OH^\u2212]= 1.0\u00d710^{- 9} M ."

"[H^+]=\frac{10^{-14}}{10^{-9}}=10^{-5 }M"


2) [OH-] = "2.6\times10^{-2}"

"[H^+]=\frac{10^{-14} }{2.6\times10^{-2}}=3.84\times10^{-13}M"


3) "[OH^-]=""7\times10^{-12}M"

"[H^+]=\frac{10^{-14}}{7\times10^{-12}}=1.43\times10^{-3}"